# Look Twice Before You Answer: Memory-Space Visual Retracing for Hallucination Mitigation in Multimodal Large Language Models

### v1

@@ 4 Experiments
This section details the evaluation of our MemVR across three MLLMs on seven benchmarks.
@@ 4.1 Experiment Setup
Datasets and Metrics. To rigorously assess the effectiveness of our proposed method, we conduct a comprehensive set of experiments across two benchmarks specifically designed to evaluate hallucination mitigation and five general-purpose benchmarks to gauge the general performance:
  • Hallucination benchmarks: Polling-based Object Probing Evaluation (POPE) (Li et al., 2023c), and Caption Hallucination Assessment with Image Relevance (CHAIR) (Rohrbach et al., 2018);
  • General-purpose benchmarks: VizWiz-VQA (Gurari et al., 2018), MLLM Comprehensive Evaluation (MME) (Fu et al., 2023), Multimodal Benchmark (MMBench) (Liu et al., 2023b), Multimodal Veterinarian (MM-Vet) (Yu et al., 2024c), LLaVA-Bench (in-the-wild) (Liu et al., 2024).
More detailed information on these various benchmarks can be obtained from the Appendix C.1.
Backbones and Baselines. To evaluate our method, we utilize three well-known MLLMs: LLaVA-1.5 (Liu et al., 2024), Qwen-VL (Bai et al., 2023), and GLM4V (Wang et al., 2023). Further, We compare our methods with classic training-tree SOTA methods designed to mitigate object hallucination, including visual contrastive decoding SOTA VCD Leng et al. (2024), OPERA (Huang et al., 2024a) based on overconfidence penalty and hindsight allocation. As Dola (Chuang et al., 2023) is layer-wise contrastive decoding for LLMs and performs poorly in MLLMs, it will not be shown in the experiment. Experimental results are obtained and benchmarked using unified implementation.
Implementation Details. Greedy search is used as the default decoding strategy in MemVR for all benchmarks. For benchmarks, annotation questions are adapted to MLLM templates. For POPE, COCO, A-OKVQA, and GQA are used, while MMBench_DEV_EN is used for MMBench. MM-Vet is assessed using MM-Vet Online Evaluator, and gpt4-1106-preview is used for LLaVA-Bench. CHAIR uses images from COCO Val2014 with the query ”Please describe this image in detail”. In MemVR, do_sample=False, temperature=0, threshold=0.75, beam=1. All settings of the compared method follow the default configurations from the original papers. More details are in Appendix C.1.
Figure 5: Results on MMBench. MemVR enhances comprehensive performance on diverse tasks.
Strategy Static-3 Static-7 Static-11 Static-15 Static-19 Static-23 Static-27 Static-31 Dynamic
Preception 1508.08 1500.48 1526.40 1529.07 1513.96 1500.29 1510.47 1510.34 1512.80
Cognition 364.64 347.14 362.86 352.14 350.36 357.86 355.71 346.43 383.92
Total Score 1872.72 1847.62 1889.26 1881.22 1864.32 1858.15 1866.18 1856.77 1896.72
Table 4: Results of MemVR-static and MemVR-dynamic on MME. Static-# indicates fixation on layer # for visual retracing. MemVR-dynamic achieves optimal performance improvements.
4.2 Quantitative Results
Key Finding: MemVR consistently outperforms the baselines in mitigating hallucinations and improving overall accuracy across various scenarios.
Figure 6: Comparison of different injection ratios $\alpha$ on cognition scores.
Q1: How does MemVR performance on hallucination benchmark?
We conduct POPE, CHAIR, and MME evaluations, as shown in Table 2, Table 3 and Table 5, our MemVR obviously surpasses all of the compared baselines. For the results of POPE evaluation, we observe that our proposed MemVR presents robust effects. The performance of MemVR surpasses the baseline results by large margins, i.e., average up to +7.0% accuracy and +7.5% F1 score on A-OKVQA dataset under Random, Popular, and Adversarial settings. As showcased in Table 5, our MemVR achieves up to 15.6% improvements on CHAIRI compared with vanilla LLaVA-1.5-7B. For MME subset evaluations (encompass both object-level and attribute-level hallucinations), results in Table 3 show that MemVR achieves a uniform improvement in handling object-level and attribute-level hallucinations, as well as commonsense reasoning. Existence, Count, and Color scores all achieve dramatic improvements (Existence score in Qwen-VL up 30). On the contrary, Position scores are relatively low, which suggests weak position reasoning capability in MLLMs.
Model Length CHAIRS $\downarrow$ CHAIRI $\downarrow$ Recall $\uparrow$
LLaVA-1.5 100.6 50.0  $\uparrow$0.0 15.4  $\uparrow$0.0 77.1  $\uparrow$0.0
+ VCD 100.4 48.6  $\downarrow$1.4 14.9  $\downarrow$0.5 77.3  $\uparrow$0.2
+ OPERA 98.6 47.8  $\downarrow$2.2 14.6  $\downarrow$0.8 76.8  $\downarrow$0.3
+ MemVR 99.6 46.6  $\downarrow$3.4 13.0  $\downarrow$2.4 80.8  $\uparrow$3.7
Table 5: CHAIR hallucination evaluation results of LLaVA. Small values correspond to fewer hallucinations.
Q2: How well does MemVR perform on general-purpose benchmarks?
@@ 4.2 Quantitative Results
We evaluate the performance of MemVR on general-purpose benchmarks, i.e., VizWiz, MME, MMBench, MM-Vet, and LLaVA-Bench. Appx. C summarizes the results on MMBench, highlighting MemVR’s comparative performance relative to SOTA methods. As shown in Table 6, MemVR consistently outperforms competing models. Besides, MemVR achieves a significant improvement in overall performance listed in Table 7, with an average increase of 6.1% in OCR and spatial awareness tasks, demonstrating superior generalization capabilities. These results indicate that compared with CD-based methods, MemVR excels in hallucination mitigation and delivers competitive performance on general-purpose benchmarks. More complete results are in Appendix C.
Model Convs $\uparrow$ Detail $\uparrow$ Complex $\uparrow$ All $\uparrow$
LLaVA-1.5 58.8  $\uparrow$0.0 52.1  $\uparrow$0.0 74.6  $\uparrow$0.0 63.4  $\uparrow$0.0
+ VCD 57.8  $\downarrow$1.0 50.8  $\downarrow$1.3 77.9  $\uparrow$3.3 59.1  $\downarrow$4.3
+ OPERA 59.5  $\uparrow$0.7 49.6  $\downarrow$2.5 78.6  $\uparrow$4.0 59.8  $\downarrow$3.6
+ MemVR 63.8  $\uparrow$0.5 52.6  $\uparrow$0.5 77.9  $\uparrow$3.3 64.0  $\uparrow$0.6
Table 6: LLaVA-Bench evaluation results.
Q3: How efficient is MemVR?
MemVR operates dynamically based on the uncertainty, which employs visual retracing when the uncertainty exceeds threshold $\gamma$ on the early layer. If the uncertainty remains low across all layers—indicating that the model is highly confident in its generated results—MemVR is not triggered. This mechanism ensures efficient inference without extra computational overhead. Compared with VCD and OPERA, they need inference twice or the rollback strategy leads to exponentially added overheads, our MemVR only once regular inference.
Figure 7: MME evaluation results under different mixing ratios of noise and visual features.
Q4: How does visual uncertainty affect MemVR?
To examine the impact of visual uncertainty on MemVR ’s performance, we introduced Gaussian noise (McHutchon & Rasmussen, 2011) into the extracted visual features. We gradually increased the noise level to assess how MemVR ’s performance would respond, using MME as the benchmark and LLaVA as the backbone model. As illustrated in Figure 7, both the perception and cognition scores declined as the noise step increased. At high noise, the performance fell significantly below the baseline. This demonstrates that MemVR is sensitive to the quality of visual features and is efficient at understanding shallow information.
Figure 8: Results of GLM4V-9B. MemVR enhances comprehensive performance on diverse tasks.
Method gn_kw_rec rec ocr_sp ocr ocr_sp_rec ocr_kw_rec ocr_gn_sp Total
LLaVA1.5-7B 18.1  $\uparrow$0.0 67.6  $\uparrow$0.0 17.7  $\uparrow$0.0 48.3  $\uparrow$0.0 60.0  $\uparrow$0.0 21.2  $\uparrow$0.0 10.0  $\uparrow$0.0 31.1  $\uparrow$0.0
+ VCD 19.2  $\uparrow$1.1 62.2  $\downarrow$5.4 15.8  $\downarrow$1.9 29.2  $\downarrow$20.0 42.5  $\downarrow$17.5 17.5  $\downarrow$3.7 60.0  $\uparrow$50.0 30.2  $\downarrow$1.1
+ OPERA 21.8  $\uparrow$3.7 61.9  $\downarrow$5.7 21.5  $\uparrow$3.8 51.7  $\uparrow$3.4 56.2  $\downarrow$3.8 11.2  $\downarrow$10.0 30.0  $\uparrow$1.4 32.0  $\uparrow$0.9
+ MemVR 19.5  $\uparrow$1.4 70.3  $\uparrow$2.7 23.8  $\uparrow$6.1 48.3  $\uparrow$0.0 58.8  $\downarrow$1.2 21.2  $\uparrow$0.0 30.0  $\uparrow$20.0 32.4  $\uparrow$1.3
@@
Table 7: MM-Vet evaluation results with multiple complicated multimodal tasks, where gn denotes language generation, kw means knowledge, sp denotes spatial awareness, and rec is recognition.

### v2

@@ 5.1 Experiment Setup
Datasets and Evaluation. To rigorously assess the effectiveness of our proposed MemVR, we conduct a comprehensive set of experiments across POPE benchmark (Li et al., 2023b), CHAIR (Rohrbach et al., 2018), VizWiz-VQA (Gurari et al., 2018), MME (Fu et al., 2023), MMBench (Liu et al., 2023b), MM-Vet (Yu et al., 2024b), LLaVA-Bench (in-the-wild) (Liu et al., 2024b), HallusionBench (Guan et al., 2024). More details can be found in Appendix B.1.
Implementation Details. Usually, we set $\gamma$=0.75. All settings of baseline methods follow the default configurations from the original papers. More details are in Appendix B.2.
Table 3: CHAIR evaluation results of different methods.
Methods CHAIRS $\downarrow$ CHAIRI $\downarrow$ Average $\downarrow$ Len Recall $\uparrow$
LLaVA-1.5 50.0  $\downarrow$0.0 15.4  $\downarrow$0.0 32.7  $\downarrow$0.0 100.6 77.1  $\uparrow$0.0
+ OPERA 47.8  $\downarrow$2.2 14.6  $\downarrow$0.8 31.2  $\downarrow$0.5 98.6 76.8  $\downarrow$0.3
+ VCD 48.6  $\downarrow$1.4 14.9  $\downarrow$0.5 31.8  $\downarrow$0.5 100.4 77.3  $\uparrow$0.2
+ ICD 56.2  $\uparrow$6.2 16.3  $\uparrow$0.9 36.3  $\uparrow$3.6 103.4 16.3  $\downarrow$60.
+ MemVR 46.6  $\downarrow$3.4 13.0  $\downarrow$2.4 29.8  $\downarrow$0.5 99.6 80.8  $\uparrow$3.7
Table 4: Performance evaluation on POPE benchmark. The best results are in green. We report accuracy and f1-score under three settings, e.g., Random, Popular, Adversarial, and Average, to show the robustness of the different methods directly.
Evaluation Methods  Random  Popular  Adversarial  Average
Accuracy $\uparrow$ F1-score $\uparrow$ Accuracy $\uparrow$ F1-score $\uparrow$ Accuracy $\uparrow$ F1-score $\uparrow$ Accuracy $\uparrow$ F1-score $\uparrow$
MSCOCO LLaVA-1.5-7B 83.49  $\uparrow$0.0 82.28  $\uparrow$0.0 79.98  $\uparrow$0.0 79.34  $\uparrow$0.0 76.03  $\uparrow$0.0 76.26  $\uparrow$0.0 79.83  $\uparrow$0.0 79.29  $\uparrow$0.0
OPERA (Huang et al., 2024a) 87.53  $\uparrow$4.0 86.45  $\uparrow$4.2 84.21  $\uparrow$4.2 83.50  $\uparrow$4.2 80.88  $\uparrow$4.9 80.69  $\uparrow$4.4 84.21  $\uparrow$4.4 83.55  $\uparrow$4.3
ICD (Wang et al., 2024) 84.87  $\uparrow$1.4 83.27  $\uparrow$1.0 82.93  $\uparrow$3.0 81.45  $\uparrow$2.1 81.07  $\uparrow$5.0 79.96  $\uparrow$3.7 82.96  $\uparrow$3.1 81.56  $\uparrow$2.3
VCD (Leng et al., 2024) 86.84  $\uparrow$3.4 86.83  $\uparrow$4.6 82.65  $\uparrow$2.7 83.37  $\uparrow$4.0 77.31  $\uparrow$1.3 79.28  $\uparrow$3.0 82.27  $\uparrow$2.4 83.16  $\uparrow$3.9
MemVR (ours) 88.50  $\uparrow$5.0 87.34  $\uparrow$5.0 87.10  $\uparrow$7.1 86.01  $\uparrow$6.7 85.20  $\uparrow$9.2 84.28  $\uparrow$8.0 86.93  $\uparrow$7.1 85.88  $\uparrow$6.6
A-OKVQA LLaVA-1.5-7B 83.45  $\uparrow$0.0 82.56  $\uparrow$0.0 79.90  $\uparrow$0.0 79.59  $\uparrow$0.0 74.04  $\uparrow$0.0 75.15  $\uparrow$0.0 79.13  $\uparrow$0.0 79.10  $\uparrow$0.0
OPERA (Huang et al., 2024a) 88.27  $\uparrow$4.8 87.54  $\uparrow$5.0 85.17  $\uparrow$5.3 84.74  $\uparrow$5.2 79.37  $\uparrow$5.3 79.97  $\uparrow$4.8 84.27  $\uparrow$5.1 84.08  $\uparrow$5.0
ICD (Wang et al., 2024) 85.57  $\uparrow$2.1 85.06  $\uparrow$2.5 81.93  $\uparrow$2.0 81.95  $\uparrow$2.4 77.43  $\uparrow$3.4 78.99  $\uparrow$3.8 81.64  $\uparrow$2.5 82.00  $\uparrow$2.9
VCD (Leng et al., 2024) 86.15  $\uparrow$2.7 86.34  $\uparrow$3.8 81.85  $\uparrow$2.0 82.82  $\uparrow$3.2 74.97  $\uparrow$0.9 77.73  $\uparrow$2.6 80.99  $\uparrow$1.9 82.30  $\uparrow$3.2
MemVR (ours) 91.10  $\uparrow$7.7 90.83  $\uparrow$8.3 87.33  $\uparrow$7.4 87.43  $\uparrow$7.8 80.20  $\uparrow$6.2 81.66  $\uparrow$6.5 86.21  $\uparrow$7.1 86.64  $\uparrow$7.5
GQA LLaVA-1.5-7B 83.73  $\uparrow$0.0 82.95  $\uparrow$0.0 78.17  $\uparrow$0.0 78.37  $\uparrow$0.0 75.08  $\uparrow$0.0 76.06  $\uparrow$0.0 78.99  $\uparrow$0.0 79.13  $\uparrow$0.0
OPERA (Huang et al., 2024a) 88.27  $\uparrow$4.5 87.52  $\uparrow$4.6 83.07  $\uparrow$4.9 82.93  $\uparrow$4.6 80.77  $\uparrow$5.7 81.05  $\uparrow$5.0 84.04  $\uparrow$5.1 83.83  $\uparrow$4.7
ICD (Wang et al., 2024) 84.90  $\uparrow$1.2 84.22  $\uparrow$1.3 78.37  $\uparrow$0.2 78.81  $\uparrow$0.4 75.97  $\uparrow$0.9 76.93  $\uparrow$0.9 79.75  $\uparrow$0.8 79.99  $\uparrow$0.9
VCD (Leng et al., 2024) 86.65  $\uparrow$2.9 86.99  $\uparrow$4.0 80.73  $\uparrow$2.6 82.24  $\uparrow$3.9 76.09  $\uparrow$1.0 78.78  $\uparrow$2.7 81.16  $\uparrow$2.2 82.67  $\uparrow$3.5
MemVR (ours) 89.60  $\uparrow$5.9 89.32  $\uparrow$6.4 84.63  $\uparrow$6.5 84.98  $\uparrow$6.6 81.53  $\uparrow$6.4 82.48  $\uparrow$6.4 85.25  $\uparrow$6.3 85.59  $\uparrow$6.5
Table 5: HallusionBench evaluation results of different methods.
Methods fACC $\uparrow$ qACC $\uparrow$ ${easy}$A $\uparrow$ ${hard}$A $\uparrow$ aACC $\uparrow$
LLaVA-1.5 17.9  $\uparrow$0.0 8.13  $\uparrow$0.0 36.0  $\uparrow$0.0 36.7  $\uparrow$0.0 41.5  $\uparrow$0.0
+ OPERA 16.2  $\downarrow$1.8 5.49  $\downarrow$2.6 37.6  $\uparrow$0.9 35.4  $\downarrow$1.3 41.2  $\downarrow$0.3
+ ICD 13.9  $\downarrow$4.0 8.35  $\uparrow$0.2 36.9  $\uparrow$0.2 33.5  $\downarrow$3.4 38.2  $\downarrow$3.3
+ VCD 13.9  $\downarrow$4.0 11.4  $\uparrow$3.3 33.0  $\downarrow$3.7 34.7  $\downarrow$3.0 41.1  $\downarrow$0.4
+ MemVR 17.9  $\uparrow$0.0 9.01  $\uparrow$0.9 36.9  $\uparrow$0.9 37.7  $\uparrow$1.0 42.5  $\uparrow$1.0
5.2 Results on Hallucination Benchmarks
We conduct hallucination evaluations on CHAIR, POPE, and HallusionBench with results presented in Table 3, Table 4 and Table 5. In the POPE evaluation, MemVR demonstrates robust effects. Its performance significantly exceeds baseline results, with an average accuracy increase of up to +7.0% and an F1-score increase of up to +7.5% on the A-OKVQA dataset under the Random, Popular, and Adversarial settings. MemVR clearly outperforms all compared SOTA methods. As shown in Table 3, compared with vanilla LLaVA-1.5, MemVR achieves 6.8% and 15.6% improvement on CHAIRS and CHAIRI metrics. Table 5 shows the HallusionBench evaluation results, where MemVR outperforms vanilla LLaVA-1.5 and other compared methods, achieving the best performance and consistent improvement across different metrics such as $hard$aACC and aACC.
5.3 Results on General-purpose Benchmarks
@@
We evaluate the performance of MemVR on general-purpose benchmarks, i.e., LLaVA-Bench, MM-Vet, MME, MMBench, and VizWiz. As shown in Table 6, MemVR consistently outperforms competing models on LLaVA-Bench. Besides, MemVR achieves a significant improvement in overall performance listed in Table 7, with an average increase of 6.1% in OCR and spatial awareness tasks, demonstrating superior generalization capabilities. For MME subset evaluations (covering both object-level and attribute-level hallucinations), the results in Figure 2 and Table 8 indicate that MemVR uniformly improves in handling object-level and attribute-level hallucinations, as well as commonsense reasoning. The Existence, Count, and Color scores all show significant improvements, where the total score of MME in LLaVA and Qwen-VL increases by 32 and 36, respectively. Importantly, although those comparison methods alleviate some hallucinations, they deteriorate generalizability, i.e., they lead to performance degradation on the general benchmarks. These results indicate that, compared to CD-based methods, MemVR excels at hallucination mitigation and delivers strong performance on general-purpose benchmarks. More experimental results are provided in Appendix B.
Figure 6: (Left) Comparisons of MemVR with regular LLaVA-1.5, Qwen-VL, and GLM-4V on LLaVA-Bench (in-the-wild). (Right) The comparison between MemVR and VCD on LLaVA-Bench shows that MemVR improves significantly while VCD decreases.
Table 6: LLaVA-Bench (In-the-Wild) evaluation results.
Methods Convs $\uparrow$ Detail $\uparrow$ Complex $\uparrow$ All $\uparrow$ Average $\uparrow$
LLaVA-1.5 58.8  $\uparrow$0.0 52.1  $\uparrow$0.0 74.6  $\uparrow$0.0 63.4  $\uparrow$0.0 64.8  $\uparrow$0.0
+ OPERA 59.5  $\uparrow$0.7 49.6  $\downarrow$2.5 78.6  $\uparrow$4.0 59.8  $\downarrow$3.6 64.3  $\downarrow$0.5
+ ICD 40.3  $\downarrow$18. 42.2  $\downarrow$9.9 60.3  $\downarrow$14. 49.8  $\downarrow$13. 56.9  $\downarrow$7.9
+ VCD 57.8  $\downarrow$1.0 50.8  $\downarrow$1.3 77.9  $\uparrow$3.3 59.1  $\downarrow$4.3 63.2  $\downarrow$1.6
+ MemVR 63.8  $\uparrow$5.0 52.6  $\uparrow$0.5 77.9  $\uparrow$3.3 64.0  $\uparrow$0.6 65.2  $\uparrow$0.4
@@
Table 7: MM-Vet evaluation results, where G denotes language Generation, K: Knowledge S: Spatial awareness, R: Recognition.
Methods R $\uparrow$ OCR_S $\uparrow$ OCR_K_R $\uparrow$ OCR_G_S $\uparrow$ Total $\uparrow$
LLaVA-1.5 67.6  $\uparrow$0.0 17.7  $\uparrow$0.0 21.2  $\uparrow$0.0 10.0  $\uparrow$0.0 31.1  $\uparrow$0.0
+ OPERA 61.9  $\downarrow$5.7 21.5  $\uparrow$3.8 11.2  $\downarrow$10. 30.0  $\uparrow$20. 32.0  $\uparrow$0.9
+ ICD 59.5  $\downarrow$8.1 17.7  $\uparrow$0.0 8.8  $\downarrow$12.4 40.0  $\uparrow$30. 25.9  $\downarrow$5.2
+ VCD 62.2  $\downarrow$5.4 15.8  $\downarrow$1.9 17.5  $\downarrow$3.7 60.0  $\uparrow$50. 30.2  $\downarrow$1.1
+ MemVR 70.3  $\uparrow$2.7 23.8  $\uparrow$6.1 21.2  $\uparrow$0.0 30.0  $\uparrow$20. 32.4  $\uparrow$1.3
Table 8: Performance evaluation on MME Hallucination subset, MM-Vet, Vizwiz, and MMBench. The best results are in green.
Evaluation Methods MME-Hall Object-Level Attribute-Level MM-Vet Vizwiz MMBench
Total $\uparrow$ Existence $\uparrow$ Count $\uparrow$ Position $\uparrow$ Color $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$
LLaVA-1.5 Regular 643.3  $\uparrow$0.0 190.0  $\uparrow$0.0 155.0  $\uparrow$0.0 128.3  $\uparrow$0.0 170.0  $\uparrow$0.0 31.10  $\uparrow$0.0 50.00  $\uparrow$0.0 62.80  $\uparrow$0.0
OPERA 610.0  $\downarrow$33. 195.0  $\uparrow$5.0 128.3  $\downarrow$26. 121.7  $\downarrow$6.6 165.0  $\downarrow$5.0 32.00  $\uparrow$0.9 50.76  $\uparrow$0.8 62.80  $\uparrow$0.0
ICD 583.3  $\downarrow$60. 185.0  $\downarrow$5.0 130.0  $\downarrow$25. 121.7  $\downarrow$6.6 146.7  $\downarrow$23. 25.90  $\downarrow$5.2 37.62  $\downarrow$12. 39.78  $\downarrow$23.
VCD 648.3  $\uparrow$5.0 190.0  $\uparrow$0.0 155.0  $\uparrow$0.0 133.3  $\uparrow$5.0 170.0  $\uparrow$0.0 30.20  $\downarrow$0.9 44.90  $\downarrow$5.1 54.21  $\downarrow$8.6
MemVR (ours) 648.3  $\uparrow$5.0 190.0  $\uparrow$0.0 155.0  $\uparrow$0.0 133.3  $\uparrow$5.0 170.0  $\uparrow$0.0 32.40  $\uparrow$1.3 51.50  $\uparrow$1.5 63.75  $\uparrow$0.9
Qwen-VL Regular 618.3  $\uparrow$0.0 175.0  $\uparrow$0.0 140.0  $\uparrow$0.0 123.3  $\uparrow$0.0 180.0  $\uparrow$0.0 49.00  $\uparrow$0.0 66.05  $\uparrow$0.0 56.53  $\uparrow$0.0
OPERA  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0
ICD 616.7  $\downarrow$1.7 170.0  $\downarrow$5.0 138.3  $\downarrow$1.7 148.3  $\uparrow$25. 160.0  $\downarrow$20. 31.70  $\downarrow$17. 29.37  $\downarrow$36. 13.32  $\downarrow$43.
VCD 648.3  $\uparrow$30. 175.0  $\uparrow$0.0 130.0  $\downarrow$10. 153.3  $\uparrow$30. 190.0  $\uparrow$10. 34.60  $\downarrow$14. 34.54  $\downarrow$31. 39.18  $\downarrow$17.
MemVR (ours) 638.3  $\uparrow$20. 185.0  $\uparrow$10. 145.0  $\uparrow$5.0 123.3  $\uparrow$0.0 185.0  $\uparrow$5.0 49.60  $\uparrow$0.6 66.36  $\uparrow$0.3 56.44  $\downarrow$0.1
5.4 Ablation Studies
To validate the effectiveness across the factors we introduced in our MemVR, we conducted in-depth ablation experiments as detailed in Figure 7 and Table 10. All ablation experiments are conducted on LLaVA-1.5-7B, and the performance is assessed across benchmarks.
Impact of Threshold and Injection Ratio. Fig. 7 shows the performance under different injection ratios $\alpha$ and thresholds $\gamma$, we find that setting $\gamma$ between 0.6 and 0.95 can improve performance, with the optimal threshold around 0.75. Understandably, a high threshold makes VR hard to trigger, while a low threshold would trigger VR at an earlier layer. For injection rate $\alpha$, we find a positive effect of between 5% and 35%, and a negative effect above 35%, which suggests that there is an upper limit to the supplementation of visual memory. To drop the hyperparameter $\alpha$ and derive the dynamic injection ratio, we calculate $\alpha$ by $\alpha=2(u-\gamma)$, where $u$ denotes uncertain score and this variant named MemVR, and the results are present in Table 9.
Figure 7: (Left) Results under different thresholds $\gamma$. (Right) Results under different injection ratios $\alpha$, both on MME benchmark.
Table 9: MemVR performance on different benchmarks.
Methods MME (Overall) POPE (COCO) Hallusion -Bench LLaVA -Bench POPE (GQA)
LLaVA-1.5 1864.68 83.49 41.45 64.80 83.73
+ MemVR 1896.72 88.50 42.34 65.17 89.60
+ MemVR 1894.14 88.40 42.27 65.87 89.57